OpenStudy (zaxoanl):

Find the equation of the line that passes through (-4,5,2) and is parallel to both the xz-plane and yz-plane. i got y-5=0 y=5 x+4=0 x=-4

OpenStudy (zaxoanl):

is (x+4)+(y-5)+(z-2)= 0 the equation?

OpenStudy (eliesaab):

No, the equation in parametric form is \[ x=-4\\ y=5\\ z=? \] Can you finish it?

OpenStudy (zaxoanl):

@eliesaab so x=-4 y=5 and z=2 is the answer to the entire question?

OpenStudy (zaxoanl):

not sure if that answer the entire question because it seems like it has 2 part

OpenStudy (eliesaab):

No the equation is x =-4 y=5 and z=t where \( -\infty < t < \infty\)

OpenStudy (eliesaab):

x=-4 y=5 and z=2 is only one point and not a whole line

OpenStudy (zaxoanl):

oh ok if the says xy and zy plane, would it be x=-4 y=t and z=2?

OpenStudy (sshayer):

x+4=y-5,z=2